OpenStudy (anonymous):

What is the length of the hypotenuse of a right triangle whose legs have lengths of 5 and 12?

OpenStudy (anonymous):

\[a^2+b^2=c^2\] where a=5 b=12 c=hypotenuse

OpenStudy (cwrw238):

use pythagoras theorem

OpenStudy (anonymous):

Plug in the values given to you and solve for c